# Factor Augmented Tensor-on-Tensor Neural Networks

### v1

@@ Abstract
Keywords: Tensor regression; Tensor-on-tensor regression; Tensor factor model;
Time series forecasting; Temporal convolutional network.
1 Introduction
@@
Motivated by a wide range of industrial and scientific applications, forecasting tasks using one tensor series to predict another has become increasingly prevalent in various fields. In finance, tensor models are used to forecast future stock prices by analyzing multi-dimensional stock attributes over time (Tran et al.,, 2017). In meteorology, tensor models predict future weather conditions using historical data formatted as multi-dimensional tensors (latitude, longitude, time, etc.) (Bilgin et al.,, 2021). In neuroscience, tensor models are employed to process medical imaging like MRI/fMRI scans (Wei et al.,, 2023).
Though predictive modeling of tensors has emerged rapidly over the past decade, most existing studies primarily focus on scenarios when either covariates or responses are tensors, such as scalar-on-tensor regression (Guo et al.,, 2011; Zhou et al.,, 2013; Wimalawarne et al.,, 2016; Li et al.,, 2018; Ahmed et al.,, 2020), matrix-on-tensor regression (Hoff,, 2015; Kossaifi et al.,, 2020), and tensor response regression (Li and Zhang,, 2017; Sun and Li,, 2017). Tensor-on-tensor [^1] Following the literature, we distinguish “tensor-on-tensor regression” from “tensor regression”. “Tensor regression” refers to regression models with tensor covariates and scalar/vector/matrix responses. “Tensor-on-tensor regression” refers to regression models with both tensor covariates and tensor responses. As a matter of fact, tensor-on-tensor regression encompasses tensor regression. prediction, when both covariates and responses are tensors, remains a challenging task due to the inherent complexities of tensor structures.
@@
One commonly used strategy for tensor-on-tensor modeling is to flatten the tensors into matrics (or even vectors) (Kilmer and Martin,, 2011; Choi and Vishwanathan,, 2014) so that matrix-based analytical tools become applicable. However, flattening can lead to a loss of spatial or temporal information, especially for datasets that are inherently dependent on data structures. For instance, in image processing, the relative positions of pixels carry important information about object shapes, which is lost when the image is flattened.
@@
Alternatively, there is a handful of recent works on tensor-on-tensor regression models that do not adopt flattening but deal with the tensor structures directly; see Lock, (2018); Liu et al., (2020); Gahrooei et al., (2021); Luo and Zhang, (2022). However, these methods posit a linear prediction model between covariates and responses, limiting their capacity to capture potential nonlinear relationships.
@@
Another emerging trend in tensor-on-tensor prediction involves neural networks, such as recurrent neural networks, convolutional neural networks, convolutional recurrent neural networks, and temporal convolutional networks, to name a few. Deep neural network is a powerful tool in prediction tasks owing to its ability to learn intricate patterns from complex datasets. Structured with multiple layers of interconnected neurons, neural networks excel in capturing nonlinearity in data. Nevertheless, they often operate in a black-box manner and typically require extensive computational resources.
@@
In this work, we develop a Factor Augmented Tensor-on-Tensor Neural Network (FATTNN) for forecasting a sequence of tensor responses using tensor covariates of arbitrary tensor order and data dimensions. One key innovation of our method is the integration of tensor factor models into deep neural networks for tensor-on-tensor regression. Factor models have been a widely utilized tool in matrix/vector data analysis (Bai and Ng,, 2002; Stock and Watson,, 2002; Pan and Yao,, 2008; Lam and Yao,, 2012) for understanding common dependence among multi-dimensional covariates. In recent years, researchers have advanced the methodology of factor models to the context of tensor time series (Chen et al., 2022a, ; Chen et al., 2022b, ; Han et al.,, 2020, 2024, 2022; Han and Zhang,, 2022). In our models, we borrow the strengths from tensor factor models, offering a more general approach than traditional vector factor analysis for prediction (Sen et al.,, 2019; Fan et al.,, 2017; Yu et al.,, 2022).
@@
  • FATTNN provides a comprehensive framework for the general supervised learning question of predicting tensor responses using tensor covariates. Although initially introduced under the problem settings of tensor-on-tensor time series forecasting, the proposed FATTNN framework is not limited to time series data, but can accommodate a variety of data types. With different types of factor models and choices of neural network architectures tailored to specific characteristics of the data, FATTNN can be naturally adapted to tensor-on-tensor regression for various data types.
This paper is organized as follows. Section 2 introduces some useful tensor terminology and sets up the problem of interest. Section 3 presents the methodological details of the proposed FATTNN algorithms. Section 4 includes applications to three real-world datasets to demonstrate the finite-sample performance of our proposed methods. Section 5 concludes the paper with discussions. Theoretical proofs, simulation studies, and supplemental numerical results are provided in appendices.
2 Preliminaries
@@
Notation. Before proceeding, we first set up some notation. Let $\mathcal{X}\in\mathbb{R}^{d_{1}\times\dots\times d_{K}}$ denote a $K$-th order tensor of dimension $d_{1}\times d_{2}\times\dots\times d_{K}$. Let $\mathcal{X}_{i_{1},\dots,i_{K}}$ denote the $({i_{1},\dots,i_{K}})$-th entry of the tensor $\mathcal{X}$ and $\mathcal{X}[\mathcal{I}_{1},\dots,\mathcal{I}_{K}]$ denotes the sub-tensor of $\mathcal{X}$ for $\mathcal{I}_{1}\subseteq\{1,\dots,d_{1}\},\dots,\mathcal{I}_{K}\subseteq\{1,\dots,d_{K}\}$. The matricization operation $\hbox{\rm mat}_{k}(\cdot)$ unfolds an order-$K$ tensor along mode $k$ to a matrix, say ${\cal X}\in\mathbb{R}^{d_{1}\times\dots\times d_{K}}$ to $\hbox{\rm mat}_{k}({\cal X})\in\mathbb{R}^{d_{k}\times d_{-k}}$ where $d_{k}=\prod_{j\neq k}d_{j}$ and its detailed definition is provided in the appendix. The Frobenius norm of tensor ${\cal X}$ is defined as $\|{\cal X}\|_{\rm F}=(\sum_{i_{1},i_{2},\cdots,i_{K}}{\cal X}_{i_{1},\dots,i_{K}}^{2})^{1/2}$. The Tucker rank of an order-$K$ tensor ${\cal X}$, denoted by Tucrank$({\cal X})$, is defined as a $K$-tuple $(r_{1},...,r_{K})$, where $r_{k}=\text{rank}(\hbox{\rm mat}_{k}({\cal X}))$. Any Tucker rank-$(r_{1},...,r_{K})$ tensor ${\cal X}$ admits the following Tucker decomposition (Tucker,, 1966): ${\cal X}={\cal S}\times_{1}U_{1}\times_{2}\cdots\times_{K}U_{K}$, where ${\cal S}\in\mathbb{R}^{r_{1}\times\cdots\times r_{K}}$ is the core tensor and $U_{k}=\text{SVD}_{r_{k}}(\hbox{\rm mat}_{k}({\cal X}))$ is the mode-$k$ top $r_{k}$ left singular vectors. Here, the mode-$k$ product of ${\cal X}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$ with a matrix $B\in\mathbb{R}^{d_{k}\times r_{k}}$, denoted by ${\cal X}\times_{k}B$, is a $d_{1}\times\cdots\times d_{k-1}\times r_{k}\times d_{k+1}\times\cdots\times d_{K}$-dimensional tensor, and its detailed definition is provided in Section 3 and Appendix A. The following abbreviations are used to denote the tensor-matrix product along multiple modes: ${\cal X}\times_{k=1}^{K}U_{k}={\cal X}\times_{1}U_{1}\times_{2}\cdots\times_{K}U_{K}$ and ${\cal X}\times_{\ell\neq k}U_{\ell}={\cal X}\times_{1}U_{1}\times_{2}\cdots\times_{k-1}U_{k-1}\times_{k+1}U_{k+1}\times_{k+2}\cdots\times_{K}U_{K}$. For any two matrices $A\in\mathbb{R}^{m_{1}\times r_{1}},B\in\mathbb{R}^{m_{2}\times r_{2}}$, denote the Kronecker product $\odot$ as $A\odot B\in\mathbb{R}^{m_{1}m_{2}\times r_{1}r_{2}}$. For any two tensors ${\cal A}\in\mathbb{R}^{m_{1}\times m_{2}\times\cdots\times m_{K}},{\cal B}\in\mathbb{R}^{r_{1}\times r_{2}\times\cdots\times r_{N}}$, denote the tensor product $\otimes$ as ${\cal A}\otimes{\cal B}\in\mathbb{R}^{m_{1}\times\cdots\times m_{K}\times r_{1}\times\cdots\times r_{N}}$, such that $({\cal A}\otimes{\cal B})_{i_{1},...,i_{K},j_{1},...,j_{N}}=({\cal A})_{i_{1},...,i_{K}}({\cal B})_{j_{1},...,j_{N}}.$
@@
3 Methodology: Factor Augmented Tensor-on-Tensor time series regression
In this section, we introduce our Factor Augmented Tensor-on-Tensor Neural Network, dubbed as FATTNN. Our work explores a broad category of challenges known as tensor-on-tensor regression, which seeks to elucidate the relationships between covariates and responses that may manifest as scalars, vectors, matrices, or higher-order tensors. As illustrated in Figure 1, our method consists of two components. First, factor tensor features are extracted by low-rank tensor factorization. Second, we develop a Tensor-on-Tensor Neural Network based on Temporal Convolutional Network (TCN) (Bai et al.,, 2018). In Section 3.1, we present a tensor factor model. This model can capture global patterns among the observed time series covariate tensors, by representing each of the original covariate tensor ${\cal X}_{t}$ as a multilinear combination of $r_{1}\times\cdots\times r_{K}$ basis across each tensor mode $k$ plus noise, where $r_{k}\ll d_{k}$. It offers a more general approach than the traditional vector factor analysis for prediction (Sen et al.,, 2019; Fan et al.,, 2017; Luo et al.,, 2022; Yu et al.,, 2022; Fan and Gu,, 2023), and preserves the tensor structures of features. It also provides dimension reduction, which significantly reduces computational complexity. In Section 3.2, we describe how the output from the global tensor factor model can be used as an input covariate for a TCN to predict future response tensors.
Figure 1: A graphical illustration of the proposed FATTNN.
The input is the observed time series $\{({\cal X}_{t},{\cal Y}_{t})\}_{t=1}^{n}$ and the new covariates $\{{\cal X}_{t}\}_{t=n+1}^{n+m}$.
The output is the forecasted future tensor responses, denoted by $\{\widehat{\cal Y}_{t}\}_{t=n+1}^{n+m}$
3.1 Tensor Factor Model (TFM)
@@
where ${\cal E}_{t}$ is idiosyncratic noise of the tensor ${\cal X}_{t}$, and $t$ is a time point. Here the $k$-mode product of $\mathcal{X}\in\mathbb{R}^{d_{1}\times d_{2}\times\ldots\times d_{K}}$ with a matrix $U\in\mathbb{R}^{d_{k}^{\prime}\times d_{k}}$, denoted as $\mathcal{X}\times_{k}U$, is an order $K$-tensor of size $d_{1}\times\ldots\times d_{k-1}\times d_{k}^{\prime}\times d_{k+1}\times\ldots\times d_{K}$ such that $(\mathcal{X}\times_{k}U)_{i_{1},\ldots,i_{k-1},j,i_{k+1},\ldots,i_{K}}=\sum_{i_{k}=1}^{d_{k}}\mathcal{X}_{i_{1},i_{2},\ldots,i_{K}}U_{j,i_{k}}.$ If $K=2$, the matrix factor model can be written as ${\cal X}_{t}=A_{1}{\cal F}_{t}A_{2}^{\top}+{\cal E}_{t}$. The latent core tensor ${\cal F}_{t}$, which typically encapsulates the most critical information, generally possesses dimensions significantly smaller than those of ${\cal X}_{t}$. Thus, the covariate tensor ${\cal X}_{t}$ can be thought of to be comprised of a basis tensor features ${\cal F}_{t}$ that capture the global patterns in the whole data-set, and all the original covariates are multilinear combinations of these basis tensor features. In the following subsection, we will discuss how a TCN can be utilized to leverage the temporal structure of the training data.
@@
The estimation methods in (2) can be further enhanced through an iterative refinement algorithm, detailed in the appendix. The factor tensors in the training data thus can be estimated by
^$\widehat{\cal F}_{t}={\cal X}_{t}\times_{k=1}^{K}\widehat{A}_{k}^{\top},\quad\text{for }1\leq t\leq n.$
@@
$\|\widehat{A}_{k}(\widehat{A}_{k}^{\top}\widehat{A}_{k})^{-1}\widehat{A}_{k}^{\top}-A_{k}(A_{k}^{\top}A_{k})^{-1}A_{k}\|_{2}=O_{\mathbb{P}}\left(\frac{\max_{k}\sqrt{d_{k}}}{\lambda\sqrt{n}}+\frac{\max_{k}\sqrt{d_{k}}}{\lambda^{2}\sqrt{n}}\right),\qquad 1\leq k\leq K.$
3.2 Tensor-on-Tensor Neural Network based TCN (TTNN)
If we are equipped with a TCN that identifies the temporal patterns in the training dataset $({\cal X}^{(tr)},{\cal Y}^{(tr)})$, we can use this model to enhance the temporal structures in $({\cal F}^{(tr)},{\cal Y}^{(tr)})$. Thus, a TCN is applied to the input sequence of estimated factor tensors ^$\widehat{\cal F}_{1},\ldots,\widehat{\cal F}_{t}$ along with the corresponding output sequence of response tensors ${\cal Y}_{1},\ldots,{\cal Y}_{t}$ to encapsulate the temporal dynamics. Let $\eta(\cdot)$ represent this network. The whole temporal network $\eta(\cdot)$ can be trained using the low-rank factor tensors ^$\widehat{\cal F}_{1},\ldots,\widehat{\cal F}_{n}$ derived from the training dataset.
The trained temporal network $\eta(\cdot)$ can be effectively utilized for multi-step look-ahead prediction in a standard approach. The factor tensors for future time steps can be estimated using the following formula:
^$\widehat{\cal F}_{n+h}={\cal X}_{n+h}\times_{k=1}^{K}\widehat{A}_{k}^{\top},$
where $h>0$. Using the historical data points of factor tensors ^$\widehat{\cal F}_{t},1\leq t\leq n+h-1$, the prediction for the subsequent time step, ${\cal Y}_{n+h}$, is generated by
$\widehat{\cal Y}_{n+h}=\eta(\widehat{\cal F}_{1},...,\widehat{\cal F}_{n+h}).$
Our hybrid forecasting model incorporates a TCN that inputs past data points from the original raw response time series and the estimated factor tensors, enhancing prediction accuracy. The pseudo-code for our model is outlined in Algorithm 1.
@@ 3.2 Tensor-on-Tensor Neural Network based TCN (TTNN)
3:  Compute the factor tensor ^$\widehat{\cal F}_{t}$ for $t=1,\ldots,n$ by TIPUP (2) or its iterative version in Algorithm S.1.
@@
4 Numerical Results
In this section, we investigate the finite-sample performance of our proposed FATTNN methods via simulation studies and real-world applications. Specifically, we implement the proposed FATTNN [^2] Our code for implementing FATTNN is available in the supplement. with TIPUP algorithm to estimate the factor structures. For benchmark methods [^3] Multiway is implemented using the code provided by https://github.com/lockEF/MultiwayRegression, and TCN is implemented using the provided by https://github.com/locuslab/TCN . , we consider one traditional statistical model - the multiway tensor-on-tensor regression (Lock,, 2018) (denoted by Multiway), and one conventional deep learning approach – the temporal convolutional network (Bai et al.,, 2018) of $\mathcal{Y}$ regressed on $\mathcal{X}$ (denoted by TCN). Due to space limitations, simulation results are presented in Appendix 4.1.

### v2

@@ Abstract
Introduction
@@
Motivated by a wide range of industrial and scientific applications, forecasting tasks using one tensor series to predict another have become increasingly prevalent in various fields. In finance, tensor models are used to forecast future stock prices by analyzing multi-dimensional stock attributes over time (Tran et al. 2017). In meteorology, tensor models predict future weather conditions using historical data formatted as multi-dimensional tensors (latitude, longitude, time, etc.) (Bilgin et al. 2021). In neuroscience, tensor models are employed to process medical imaging like MRI/fMRI scans (Wei et al. 2023). Though predictive modeling of tensors has emerged rapidly over the past decade, most existing studies primarily focus on scenarios when either covariates or responses are tensors, such as scalar-on-tensor regression (Guo, Kotsia, and Patras 2011; Zhou, Li, and Zhu 2013; Wimalawarne, Tomioka, and Sugiyama 2016; Li et al. 2018; Ahmed, Raja, and Bajwa 2020), matrix-on-tensor regression (Hoff 2015; Kossaifi et al. 2020), and tensor-on-vector regression (Li and Zhang 2017; Sun and Li 2017). Tensor-on-tensor [^1] Following the literature, we distinguish “tensor-on-tensor regression” from “tensor regression”. “Tensor regression” refers to regression models with tensor covariates and scalar responses. “Tensor-on-tensor regression” refers to regression models with both tensor covariates and tensor responses. As a matter of fact, tensor-on-tensor regression encompasses tensor regression. prediction, when both covariates and responses are tensors, remains a challenging task due to the inherent complexities of tensor structures.
@@
One commonly used strategy for tensor-on-tensor modeling is to flatten the tensors into matrices (or even vectors) (Kilmer and Martin 2011; Choi and Vishwanathan 2014) so that matrix-based analytical tools become applicable. However, flattening can lead to a loss of spatial or temporal information, especially for datasets that are inherently dependent on data structures. For instance, in image processing, the relative positions of pixels carry important information about object shapes, which is lost when the image is flattened.
@@
Alternatively, there is a handful of recent works on tensor-on-tensor regression models that do not adopt flattening but deal with the tensor structures directly (Lock 2018; Liu, Liu, and Zhu 2020; Gahrooei et al. 2021; Luo and Zhang 2024). However, these methods posit a linear prediction model between covariates and responses, limiting their capacity to capture potential nonlinear relationships.
@@
Another emerging trend in tensor-on-tensor prediction involves neural networks, such as recurrent neural networks (RNN), convolutional neural networks (CNN), recurrent convolutional neural networks (RCNN), and temporal convolutional networks (TCN), to name a few. Deep neural network is a powerful tool in prediction tasks owing to its ability to learn intricate patterns from complex datasets. Structured with multiple layers of interconnected neurons, neural networks excel in capturing nonlinearity in data. Nevertheless, they often operate in a black-box manner and typically require extensive computational resources.
@@
In this work, we develop a Factor Augmented Tensor-on-Tensor Neural Network (FATTNN) for forecasting a sequence of tensor responses using tensor covariates of arbitrary tensor order and data dimensions. One key innovation of our method is the integration of tensor factor models into deep neural networks for tensor-on-tensor regression. Factor models have been a widely utilized tool in matrix/vector data analysis (Bai and Ng 2002; Stock and Watson 2002; Pan and Yao 2008; Lam and Yao 2012) for understanding common dependence among multi-dimensional covariates. In recent years, researchers have advanced the methodology of factor models to the context of tensor time series (Chen et al. 2022; Chen, Yang, and Zhang 2022; Han et al. 2024a, b; Han, Chen, and Zhang 2022; Han and Zhang 2023; Chen, Han, and Yu 2024; Yu et al. 2024). In our models, we borrow the strengths from tensor factor models, offering a more general approach than traditional vector factor analysis for prediction (Sen, Yu, and Dhillon 2019; Fan, Xue, and Yao 2017; Yu, Yao, and Xue 2022).
@@
  • FATTNN provides a comprehensive framework for the general supervised learning question of predicting tensor responses using tensor covariates. Although initially introduced under the problem settings of tensor-on-tensor time series forecasting, the proposed framework is not limited to time series data, but can accommodate a variety of data types. With different types of factor models and choices of neural network architectures tailored to specific characteristics of the data, FATTNN can be naturally adapted to tensor-on-tensor regression for various data types.
Related Work
Neural networks are powerful tools for processing complex tensor data. Early works involve RNN and its variations (e.g., Long Short-Term Memory (LSTM) networks) to accommodate the temporal dependence, and CNN to capture the temporal and spatial relationships underlying the data structures. Recently, more advanced architectures have been proposed, integrating the strengths of traditional models to capture the increasingly complex data dependence in a more powerful and efficient manner. Notable examples include convolutional LSTM (ConvLSTM) (Shi et al. 2015), recurrent CNN (RCNN) (Liang and Hu 2015), and temporal convolutional networks (TCN)(Bai, Kolter, and Koltun 2018).
Another popular thread is to incorporate a tensor decomposition procedure in the model training to reduce computational complexity and potentially enhance interpretability. Recent examples include tensorizing neural networks (Novikov et al. 2015), convolutional tensor-train LSTM (Conv-TT-LSTM) (Su et al. 2020), tensor regression networks (Kossaifi et al. 2020) and its efficient variant using tensor dropout (Kolbeinsson et al. 2021). Methods may vary depending on how the tensor decomposition is conducted (Fang et al. 2022; Wang and Zhe 2022; Tao, Tanaka, and Zhao 2024).
Preliminaries
@@
Notation. Before proceeding, we first set up some notation. Let $\mathcal{X}\in\mathbb{R}^{d_{1}\times\dots\times d_{K}}$ denote a $K$-th order tensor of dimension $d_{1}\times d_{2}\times\dots\times d_{K}$. Let $\mathcal{X}_{i_{1},\dots,i_{K}}$ denote the $({i_{1},\dots,i_{K}})$-th entry of the tensor $\mathcal{X}$ and $\mathcal{X}[\mathcal{I}_{1},\dots,\mathcal{I}_{K}]$ denotes the sub-tensor of $\mathcal{X}$ for $\mathcal{I}_{1}\subseteq\{1,\dots,d_{1}\},\dots,\mathcal{I}_{K}\subseteq\{1,\dots,d_{K}\}$. The matricization operation $\hbox{\rm mat}_{k}(\cdot)$ unfolds an order-$K$ tensor along mode $k$ to a matrix, say ${\cal X}\in\mathbb{R}^{d_{1}\times\dots\times d_{K}}$ to $\hbox{\rm mat}_{k}({\cal X})\in\mathbb{R}^{d_{k}\times d_{-k}}$ where $d_{k}=\prod_{j\neq k}d_{j}$ and its detailed definition is provided in the appendix. The Frobenius norm of tensor ${\cal X}$ is defined as $\|{\cal X}\|_{\rm F}=(\sum_{i_{1},i_{2},\cdots,i_{K}}{\cal X}_{i_{1},\dots,i_{K}}^{2})^{1/2}$. The Tucker rank of an order-$K$ tensor ${\cal X}$, denoted by Tucrank$({\cal X})$, is defined as a $K$-tuple $(r_{1},...,r_{K})$, where $r_{k}=\text{rank}(\hbox{\rm mat}_{k}({\cal X}))$. Any Tucker rank-$(r_{1},...,r_{K})$ tensor ${\cal X}$ admits the following Tucker decomposition (Tucker 1966): ${\cal X}={\cal S}\times_{1}U_{1}\times_{2}\cdots\times_{K}U_{K}$, where ${\cal S}\in\mathbb{R}^{r_{1}\times\cdots\times r_{K}}$ is the core tensor and $U_{k}=\text{SVD}_{r_{k}}(\hbox{\rm mat}_{k}({\cal X}))$ is the mode-$k$ top $r_{k}$ left singular vectors. Here, the mode-$k$ product of ${\cal X}\in\mathbb{R}^{d_{1}\times\cdots\times d_{K}}$ with a matrix $B\in\mathbb{R}^{d_{k}\times r_{k}}$, denoted by ${\cal X}\times_{k}B$, is a $d_{1}\times\cdots\times d_{k-1}\times r_{k}\times d_{k+1}\times\cdots\times d_{K}$-dimensional tensor, and its detailed definition is provided in Appendix A. The following abbreviations are used to denote the tensor-matrix product along multiple modes: ${\cal X}\times_{k=1}^{K}U_{k}={\cal X}\times_{1}U_{1}\times_{2}\cdots\times_{K}U_{K}$ and ${\cal X}\times_{\ell\neq k}U_{\ell}={\cal X}\times_{1}U_{1}\times_{2}\cdots\times_{k-1}U_{k-1}\times_{k+1}U_{k+1}\times_{k+2}\cdots\times_{K}U_{K}$. For any two matrices $A\in\mathbb{R}^{m_{1}\times r_{1}},B\in\mathbb{R}^{m_{2}\times r_{2}}$, we denote the Kronecker product $\odot$ as $A\odot B\in\mathbb{R}^{m_{1}m_{2}\times r_{1}r_{2}}$. For any two tensors ${\cal A}\in\mathbb{R}^{m_{1}\times m_{2}\times\cdots\times m_{K}},{\cal B}\in\mathbb{R}^{r_{1}\times r_{2}\times\cdots\times r_{N}}$, we denote the tensor product $\otimes$ as ${\cal A}\otimes{\cal B}\in\mathbb{R}^{m_{1}\times\cdots\times m_{K}\times r_{1}\times\cdots\times r_{N}}$, such that $({\cal A}\otimes{\cal B})_{i_{1},...,i_{K},j_{1},...,j_{N}}=({\cal A})_{i_{1},...,i_{K}}({\cal B})_{j_{1},...,j_{N}}.$
@@
Methodology: Factor Augmented Tensor-on-Tensor Time Series Regression
Figure 1: A graphical illustration of the proposed FATTNN. The input is the observed time series $\{({\cal X}_{t},{\cal Y}_{t})\}_{t=1}^{n}$ and the new covariates $\{{\cal X}_{t}\}_{t=n+1}^{n+m}$. The output is the forecasted future tensor responses, denoted by $\{\widehat{\cal Y}_{t}\}_{t=n+1}^{n+m}$
In this section, we introduce our Factor Augmented Tensor-on-Tensor Neural Network, dubbed as FATTNN. Our work explores a broad category of challenges known as tensor-on-tensor regression, which seeks to elucidate the relationships between covariates and responses that may manifest as scalars, vectors, matrices, or higher-order tensors. As illustrated in Figure 1, our method consists of two components. First, factor tensor features are extracted by low-rank tensor factorization. Second, we develop a Tensor-on-Tensor Neural Network based on Temporal Convolutional Network (TCN) (Bai, Kolter, and Koltun 2018). In the first subsection, we present a tensor factor model. This model can capture global patterns among the observed time series covariate tensors, by representing each of the original covariate tensor ${\cal X}_{t}$ as a multilinear combination of $r_{1}\times\cdots\times r_{K}$ basis across each tensor mode $k$ plus noise, where $r_{k}\ll d_{k}$. It offers a more general approach than the traditional vector factor analysis for prediction (Sen, Yu, and Dhillon 2019; Fan, Xue, and Yao 2017; Luo et al. 2022; Yu, Yao, and Xue 2022; Fan and Gu 2023), and preserves the tensor structures of features. It also provides dimension reduction, which significantly reduces computational complexity. In the second subsection, we describe how the output from the global tensor factor model can be used as an input covariate for a TCN to predict future response tensors.
Tensor Factor Model (TFM)
@@
where ${\cal E}_{t}$ is idiosyncratic noise of the tensor ${\cal X}_{t}$, and $t$ is a time point. Here the $k$-mode product of $\mathcal{X}\in\mathbb{R}^{d_{1}\times d_{2}\times\ldots\times d_{K}}$ with a matrix $U\in\mathbb{R}^{d_{k}^{\prime}\times d_{k}}$, denoted as $\mathcal{X}\times_{k}U$, is an order $K$-tensor of size $d_{1}\times\ldots\times d_{k-1}\times d_{k}^{\prime}\times d_{k+1}\times\ldots\times d_{K}$ such that $(\mathcal{X}\times_{k}U)_{i_{1},\ldots,i_{k-1},j,i_{k+1},\ldots,i_{K}}=\sum_{i_{k}=1}^{d_{k}}\mathcal{X}_{i_{1},i_{2},\ldots,i_{K}}U_{j,i_{k}}.$ If $K=2$, the matrix factor model can be written as ${\cal X}_{t}=A_{1}{\cal F}_{t}A_{2}^{\top}+{\cal E}_{t}$. The latent core tensor ${\cal F}_{t}$, which typically encapsulates the most critical information, generally possesses dimensions significantly smaller than those of ${\cal X}_{t}$. Thus, the covariate tensor ${\cal X}_{t}$ can be thought of to be comprised of a basis tensor features ${\cal F}_{t}$ that capture the global patterns in the whole data-set, and all the original covariates are multilinear combinations of these basis tensor features. In the next subsection, we will discuss how a TCN can be utilized to leverage the temporal structure of the training data.
@@
The estimation methods in (2) can be further enhanced through an iterative refinement algorithm, detailed in the appendix. The factor tensors in the training data thus can be estimated by ^$\widehat{\cal F}_{t}={\cal X}_{t}\times_{k=1}^{K}\widehat{A}_{k}^{\top},\text{for }1\leq t\leq n$.
@@
$\|\widehat{A}_{k}(\widehat{A}_{k}^{\top}\widehat{A}_{k})^{-1}\widehat{A}_{k}^{\top}-A_{k}(A_{k}^{\top}A_{k})^{-1}A_{k}\|_{2}$
$=$ $O_{\mathbb{P}}\left(\frac{\max_{k}\sqrt{d_{k}}}{\lambda\sqrt{n}}+\frac{\max_{k}\sqrt{d_{k}}}{\lambda^{2}\sqrt{n}}\right),\qquad 1\leq k\leq K.$
Tensor-on-Tensor Neural Network based TCN (TTNN)
If we are equipped with a TCN that identifies the temporal patterns in the training dataset $({\cal X}^{(tr)},{\cal Y}^{(tr)})$, we can use this model to enhance the temporal structures in $({\cal F}^{(tr)},{\cal Y}^{(tr)})$. Thus, a TCN is applied to the input sequence of estimated factor tensors ^$\widehat{\cal F}_{1},\ldots,\widehat{\cal F}_{t}$ along with the corresponding output sequence of response tensors ${\cal Y}_{1},\ldots,{\cal Y}_{t}$ to encapsulate the temporal dynamics. Let $\eta(\cdot)$ represent this network. The whole temporal network $\eta(\cdot)$ can be trained using the low-rank factor tensors ^$\widehat{\cal F}_{1},\ldots,\widehat{\cal F}_{n}$ derived from training dataset. In the literature, factor-augmented linear regression has been well studied recently in statistics, economics and machine learning (Stock and Watson 2002; Bai and Ng 2006; Bair et al. 2006; Fan, Xue, and Yao 2017; Bing et al. 2021).
The trained temporal network $\eta(\cdot)$ can be effectively utilized for multi-step look-ahead prediction in a standard approach. The factor tensors for future time steps can be estimated using ^$\widehat{\cal F}_{n+h}={\cal X}_{n+h}\times_{k=1}^{K}\widehat{A}_{k}^{\top}$, where $h>0$. Using the historical data points of factor tensors ^$\widehat{\cal F}_{t},1\leq t\leq n+h-1$, the prediction for the subsequent time step, ${\cal Y}_{n+h}$, is generated by $\widehat{\cal Y}_{n+h}=\eta(\widehat{\cal F}_{1},...,\widehat{\cal F}_{n+h})$.
Our hybrid forecasting model incorporates a TCN that inputs past data points from the original raw response time series and the estimated factor tensors, enhancing prediction accuracy. The pseudo-code for our model is outlined in Algorithm 1. Due to space limitations, the determination of the rank of the factor process is discussed in Appendix A.
@@
3:  Compute the factor tensor ^$\widehat{\cal F}_{t}$ for $t=1,\ldots,n$, by TIPUP (2) or its iterative version in Algorithm S.1.
@@
Numerical Results
In this section, we investigate the finite-sample performance of our proposed FATTNN methods via simulation studies and real-world applications. Specifically, we implement the proposed FATTNN [^2] Our code is available on Github: https://github.com/chenezandie/aaai2025-fattnn. , we consider one traditional statistical model - the multiway tensor-on-tensor regression (Lock 2018) (denoted by Multiway), and four state-of-the-art deep learning approaches, including the temporal convolutional network (Bai, Kolter, and Koltun 2018) of $\mathcal{Y}$ regressed on $\mathcal{X}$ (denoted by TCN), the long short-term memory network (Hochreiter and Schmidhuber 1997) (denoted by LSTM) , the convolutional tensor-train LSTM (Su et al. 2020) (denoted by Conv-TT-LSTM), and tensor regression layer (Kossaifi et al. 2020) (denoted by TRL).